The energy of a particular color of visible light is 4.113x10^-19. What is the wavelength of this light?
1
Expert's answer
2017-03-29T07:28:23-0400
Solution:
According to Plank’s formula:
E = hc/λ,
were E is energy; h – Plank’s constant (h = 6.626·10-34); c is speed of light (c = 2.998·108 m/s); λ – wavelength.
That’s why wavelength equals:
λ = hc/E = (6.626·〖10〗^(-34)∙2.998∙〖10〗^8)/(4.113·〖10〗^(-19) ) = 4.830·10-7 (m)

Answer: λ = 4.830·10-7 m.